# Canonical Form and Finite Blocklength Bounds for Stabilizer Codes

### v1

@@ 1 Introduction
These applications are not considered in detail here; the argument for the encoding circuit is similar to [17, Section III], where a decomposition of CSS parity check matrices is used to derive the encoding circuit and the algorithm for generation of uniformly random stabilizer parity check matrices of given size and rank is similar to the algorithm in [5, Section III] for the generation of uniformly random Clifford group elements.
@@
Theorem 7 below gives a finite blocklength refinement of the hashing bound, applicable to arbitrary $n$-qubit Pauli noise. Then, Theorems 8 and 9 show how the general bound can be computed efficiently in the important special cases of $n$ independent qubit erasure channels and $n$ independent qubit depolarizing channels.
@@ 2.2 The Pauli group
[…]
@@
is a surjective group homomorphism with kernel {±$\{\pm I,\pm iI\}$. $\Theta_{n}$ sends $Q_{2n,i}$ to $e_{2n,i}$. Additinally, the identity
@@ 2.6 Preparation and measurement
Consider now the $n$-qubit Pauli channel specified by (17). Suppose the $n$ qubits are divided in two registers: the first $m$ qubits, and the remaining $k=n-m$ qubits. Suppose further that the first $m$ qubits are prepared in the zero state, then the channel is applied, then the first $m$ qubits are measured in the computational basis. These operations transform the $n$-qubit Pauli channel into a $k$-qubit Pauli channel acting on the last $k$ qubits. The channel on the last $k$ qubits transforms $k$-qubit states as follows:
@@ 2.9 Optimal stabilizer codes and coset decoding
For $r=k/n$, let $\varepsilon_{coset}\left(p_{UV},r\right)$ be the lowest error probability of a pair $C,D$ that uses $m=n-k$ bits of syndrome.
@@ 2.10 Guessing the error from the syndrome
For $r=k/n$, let $\varepsilon_{errorguess}\left(p_{UV},r\right)$ be the lowest probability of guessing $U$ incorrectly of a pair $C,D$ that uses $m=n-k$ bits of syndrome.
@@ 3.1 A family of groups from transitive sets of pairs
Some examples
The lower triangular group
is $\mathcal{L}_{n}\left(Pairs\left(n\right)\right)$.
Matrices with off-diagonal entries only in certain rows and columns:
take subsets $R,C$ of $[n]$. Then, $T=Pairs\left(n\right)\cap(R\times C)$ is transitive. $\mathcal{L}_{n}\left(T\right)$ is a group of lower triangular matrices whose non-zero off-diagonal entries appear only in rows in $R$ and columns in $C$.
The inversions and non-inversions of a permutation:
let $\pi$ be a permutation of $[n]$. Let
@@ 3.4.1 The Clifford Gaussian moves
    For $i\leq n<j$, $W_{2n,i,j}=W_{2n,2n+1-j,2n+1-i}$ is a CZ gate conjugated by Hadamard gates on qubits $i,2n+1-j$.
@@ 3.4.2 The symplectic Gaussian moves
$S_{2n,i,j}=\Psi_{n}(W_{2n,i,j})\\ =\begin{cases}I_{2n}+e_{2n,i}e_{2n,j}^{\top}&\text{if }i+j=2n+1\\ I_{2n}+e_{2n,i}e_{2n,j}^{\top}+e_{2n,2n+1-j}e_{2n,2n+1-i}^{\top}&\text{otherwise}\end{cases}$ (41)
@@ 3.4.3 Clearing entire rows and columns
With classical Gaussian moves:
Take $i\in[n]$. Take a vector
@@
that does not have an $i$-th component (i.e. $v_{i}=0$). Let
@@
With symplectic Gaussian moves:
Take $i\in[2n]$. Take a vector
@@
that does not have an $i$-th component (i.e. $v_{i}=0$). Let
@@
    For fixed $i$ the symplectic Gaussian moves $\{S_{2n,j,i}:j\in[2n]\backslash\{i\}\}$ commute.
@@
    $S_{2n,v,i}$ is supported only on the main diagonal, column $i$ and row $2n+1-i$.
@@ 3.4.4 An identity reveals the generators and canonical form of the groups $\mathcal{B}_{2n}\left(T\right)$
Since $e_{2n,k}^{\top}B_{k-1}=e_{2n,k}^{\top}$, the second term is supported only on the $k$-th column. The third and fourth terms are supported only on row $2n+1-k$. Thus, $B_{k}-B_{k-1}$ is supported only on column $k$ and row $2n+1-k$. Moreover, column $k$ of $B_{k-1}$ is $e_{2n,k}+\mathcal{V}_{n}(B)e_{n,k}$, so column $k$ of $B_{k}$ is just $e_{2n,k}$. Finally, row $2n+1-k$ of $B_{k}$ is determined by column $k$ via the symplectic constraint:
@@ 4.1 Gaussian elimination produces a canonical form for matrices of a given size
If $A$ is partially but not fully reduced, then extend $(\alpha,\beta)$ to $(\alpha^{\prime},\beta^{\prime})$ so that $\alpha^{\prime}(r+1),\beta^{\prime}(r+1)$ is the position of the next pivot in the left and down order. Then, transform $A$:
@@
to a matrix $A^{\prime}$ that is $(\alpha^{\prime},\beta^{\prime})$-partially simplified.
Now, observe that on the left, only $G_{m,i,\alpha^{\prime}(r+1)}$ with $i>\alpha^{\prime}(r+1)$ are used. Similarly, on the right, only $G_{n,\beta^{\prime}(r+1),j}$ with $j<\beta^{\prime}(r+1)$ are used. However, on the right, there is an additional restriction: since columns $\beta(1),\dots,\beta(r)$ of $A$ are already simplified, $G_{n,\beta^{\prime}(r+1),j}$ with $j\in Im(\beta)$ are not used. [^1] If the search for pivots proceeded along columns rather than along rows, i.e. in the order last column top to bottom, then second last column top to bottom, etc., the additional restriction would be on the left rather than on the right.
@@
$T^{Right}\left(\beta\right)=\left\{(i,j):i\in Im(\beta),j<i,j\notin\left\{\beta(1),\dots,\beta\left(\beta^{-1}(i)-1\right)\right\}\right\}\\ \subset Pairs\left(n\right)$ (56)
Note that these sets are transitive and behave naturally under extensions: for $\alpha^{\prime}$ an extension of $\alpha$, $T^{Left}\left(\alpha\right)\subset T^{Left}\left(\alpha^{\prime}\right)$, and for $\beta^{\prime}$ and extension of $\beta$, $T^{Right}\left(\beta\right)\subset T^{Right}\left(\beta^{\prime}\right)$.
@@
$(r,\alpha,\beta,L,R)$ can be computed from $A$ in time $O(mnr)$ by Gaussian elimination. [^2] Assuming $L,R$ are output as lists of off-diagonal non-zero positions; otherwise, outputting the full $L,R$ would take time $\max(m,n)^{2}$.
@@
Take any $k\in[m],l\in[n]$, and divide the matrices into blocks with $(k,m-k)$ and $(l,n-l)$ rows/columns:
@@
Case 1: $i\leq j$.
Then, $e_{n,i}^{\top}Re_{n,j}=e_{n,i}^{\top}e_{n,j}$ because $R$ is lower triangular.
Case 2: $i>j$, $i\notin Im(\beta)$.
Then, $e_{n,i}^{\top}Re_{n,j}=0$ because $R\in\mathcal{L}_{n}\left(T^{Right}\left(\beta\right)\right)$.
Case 3: $i>j$, $\exists k:i=\beta(k)$, $j\notin Im(\beta)$.
Then,
@@
Case 4: $i>j$, $\exists k,l:i=\beta(k),j=\beta(l),l<k$.
Then, $e_{n,i}^{\top}Re_{n,j}=0$ because $R\in\mathcal{L}_{n}\left(T^{Right}\left(\beta\right)\right)$.
Case 5: $i>j$, $\exists k,l:i=\beta(k),j=\beta(l),k<l$.
Then,
@@
Let $\alpha:[r]\rightarrow[m]$ be increasing, and let $L$ be in $\mathcal{L}_{m}\left(T^{Right}\left(\alpha\right)\right)$. If there exist $n$ and injective $\beta:[r]\rightarrow[n]$ such that $L\Pi(\alpha,\beta)=\Pi(\alpha,\beta)$, then $L=I$.
@@ 4.2 Canonical form for $m\times 2n$ stabilizer parity check matrices
$T^{moves}\left(\beta\right)=$
$\left\{(i,j):i\in Im(\beta),j<i,q_{n}(j)\notin\left\{q_{n}(\beta(1)),\dots,q_{n}(\beta(\beta^{-1}(i)-1))\right\}\right\}$ (66)
@@
Case 1: $(i,j),(j,k)\in T^{moves}\left(\beta\right)$.
Then, $i,j\in Im(\beta)$ with $\beta^{-1}(j)>\beta^{-1}(i)$. Moreover, $i>j>k$ and
@@
Case 2: $(i,j),(j,k)\in T^{rev}\left(\beta\right)$.
Then, $(2n+1-k,2n+1-j),(2n+1-j,2n+1-i)\in T^{moves}\left(\beta\right)$, so $(2n+1-k,2n+1-i)\in T^{moves}\left(\beta\right)$, so $(i,k)\in T^{rev}\left(\beta\right)$.
Case 3: $(i,j)\in T^{moves}\left(\beta\right),(j,k)\in T^{rev}\left(\beta\right)$.
Then $i>j>k$ and $i,2n+1-k\in Im(\beta)$. Consider subcases:
@@
Case 4: $(i,j)\in T^{rev}\left(\beta\right),(j,k)\in T^{moves}\left(\beta\right)$.
Then, $j,2n+1-j\in Im(\beta)$. This is a contradiction, because $\beta$ is qubit injective. This case does not occur.
@@
Parts 2 and 3:
Since $A$ is $(\alpha,\beta)$ reduced, and since $A\Omega_{2n}A^{\top}=0$, the following holds for the columns of $A$:
@@
Part 5:
Computation of $(\alpha^{\prime},\beta^{\prime},u,v)$ can clearly be done in time $O(mn)$. For the computation of $A^{\prime}$, use the sparsity of $G_{m,u,\alpha^{\prime}(r+1)},S_{2n,\beta^{\prime}(r+1),v^{\top}}$. For any $x\in\mathbb{F}_{2}^{m}$, $G_{m,u,\alpha^{\prime}(r+1)}x$ can be computed in time $O(m)$; since $A$ has $2n$ columns, this gives time $O(mn)$ to compute $G_{m,u,\alpha^{\prime}(r+1)}A$. Similarly, for any $y\in\mathbb{F}_{2}^{2n}$, $y^{\top}S_{2n,\beta^{\prime}(r+1),v^{\top}}$ can be computed in time $O(n)$; for $m$ rows, this gives time $O(mn)$ to compute $G_{m,u,\alpha^{\prime}(r+1)}AS_{2n,\beta^{\prime}(r+1),v^{\top}}$. ∎
@@
Case 1: $(i,j)\in T^{moves}\left(\beta\right)$.
Then, $i=\beta(k)$ for some $k\in[r]$ and
@@
Case 2: $(i,j)\in T^{rev}\left(\beta\right)$.
Then, $(2n+1-j,2n+1-i)\in T^{moves}\left(\beta\right)$ and $2n+1-j=\beta(k)$ for some $k\in[r]$. Apply the identity $R^{\top}=\Omega_{2n}R^{-1}\Omega_{2n}$, which holds for any element of $\mathcal{SP}_{2n}$:
@@
    Moreover, the algorithm runs in time $O(mnr)$.
@@
Part 1 follows by repeated application of lemma 8.
Part 2:

### v2

@@ 1 Introduction
These applications are not considered in detail here. The argument for the encoding circuit is similar to [17, Section III], where a decomposition of CSS parity check matrices is used to derive the encoding circuit. The algorithm for generation of uniformly random stabilizer parity check matrices of given size and rank is similar to the algorithm in [5, Section III] for the generation of uniformly random Clifford group elements.
@@
Theorem 7 below gives a finite blocklength refinement of the hashing bound, applicable to arbitrary $n$-qubit Pauli noise. Then, Theorems 8 and 9 show how the general bound can be computed efficiently in the important special cases of $n$ independent qubit erasure channels and $n$ independent qubit depolarizing channels. Theorem 10 shows that the bounds can be computed in polynomial time in the case of $n$ independent identical Pauli channels.
@@ 2.2 The Pauli group
The operation $\left\langle,\right\rangle$ has the following properties
@@
is a surjective group homomorphism with kernel $\{\pm I,\pm iI\}$. $\Theta_{n}$ sends $Q_{2n,i}$ to $e_{2n,i}$. Additinally, the identity
@@ 2.6 Preparation and measurement
Consider now the $n$-qubit Pauli channel specified by (17). Suppose the $n$ qubits are divided in two registers: the first $m$ qubits, and the remaining k=nmk=n-m qubits. Suppose further that the first $m$ qubits are prepared in the zero state, then the channel is applied, then the first $m$ qubits are measured in the computational basis. These operations transform the $n$-qubit Pauli channel into a $k$-qubit Pauli channel acting on the last $k$ qubits. The channel on the last $k$ qubits transforms $k$-qubit states as follows:
@@ 2.9 Optimal stabilizer codes and coset decoding
For $r=k/n$, let $\varepsilon_{coset}\left(p_{UV},r\right)$ be the lowest error probability of a pair $C,D$ that uses m=nkm=n-k bits of syndrome.
@@ 2.10 Guessing the error from the syndrome
For $r=k/n$, let $\varepsilon_{errorguess}\left(p_{UV},r\right)$ be the lowest probability of guessing $U$ incorrectly of a pair $C,D$ that uses m=nkm=n-k bits of syndrome.
@@ 3.1 A family of groups from transitive sets of pairs
Some examples:
The lower triangular group is $\mathcal{L}_{n}\left(Pairs\left(n\right)\right)$.
Matrices with off-diagonal entries only in certain rows and columns: take subsets $R,C$ of $[n]$. Then, $T=Pairs\left(n\right)\cap(R\times C)$ is transitive. $\mathcal{L}_{n}\left(T\right)$ is a group of lower triangular matrices whose non-zero off-diagonal entries appear only in rows in $R$ and columns in $C$.
The inversions and non-inversions of a permutation: let $\pi$ be a permutation of $[n]$. Let
@@ 3.4.1 The Clifford Gaussian moves
    For $i\leq n<j$, $W_{2n,i,j}=W_{2n,2n+1-j,2n+1-i}$ is a CZ gate conjugated by Hadamard gates on qubits i,2n+1ji,2n+1-j.
@@ 3.4.2 The symplectic Gaussian moves
$S_{2n,i,j}=\Psi_{n}(W_{2n,i,j})=\begin{cases}I_{2n}+e_{2n,i}e_{2n,j}^{\top}&\text{if }i+j=2n+1\\ I_{2n}+e_{2n,i}e_{2n,j}^{\top}+e_{2n,2n+1-j}e_{2n,2n+1-i}^{\top}&\text{otherwise}\end{cases}$ (41)
@@ 3.4.3 Clearing entire rows and columns
With classical Gaussian moves: Take $i\in[n]$. Take a vector
@@
such that the $i$-th component is 0 (i.e. $v_{i}=0$). Let
@@
With symplectic Gaussian moves: Take $i\in[2n]$. Take a vector
@@
such that the $i$-th component is 0 (i.e. $v_{i}=0$). Let
@@
[…]
@@
    $S_{2n,v,i}$ is supported only on the main diagonal, column $i$ and row 2n+1i2n+1-i.
@@ 3.4.4 An identity reveals the generators and canonical form of the groups $\mathcal{B}_{2n}\left(T\right)$
Since $e_{2n,k}^{\top}B_{k-1}=e_{2n,k}^{\top}$, the second term is supported only on the $k$-th column. The third and fourth terms are supported only on row 2n+1k2n+1-k. Thus, $B_{k}-B_{k-1}$ is supported only on column $k$ and row 2n+1k2n+1-k. Moreover, column $k$ of $B_{k-1}$ is $e_{2n,k}+\mathcal{V}_{n}(B)e_{n,k}$, so column $k$ of $B_{k}$ is just $e_{2n,k}$. Finally, row 2n+1k2n+1-k of $B_{k}$ is determined by column $k$ via the symplectic constraint:
@@ 4.1 Gaussian elimination produces a canonical form for matrices of a given size
If $A$ is partially but not fully reduced, then extend $(\alpha,\beta)$ to $(\alpha^{\prime},\beta^{\prime})$ so that $\alpha^{\prime}(r+1),\beta^{\prime}(r+1)$ is the position of the next pivot in the left and down order. Then, transform $A$ to a matrix $A^{\prime}$ that is $(\alpha^{\prime},\beta^{\prime})$-partially reduced:
@@
Now, observe that on the left, only $G_{m,i,\alpha^{\prime}(r+1)}$ with $i>\alpha^{\prime}(r+1)$ are used. Similarly, on the right, only $G_{n,\beta^{\prime}(r+1),j}$ with $j<\beta^{\prime}(r+1)$ are used. However, on the right, there is an additional restriction: since columns $\beta(1),\dots,\beta(r)$ of $A$ are already reduced, $G_{n,\beta^{\prime}(r+1),j}$ with $j\in Im(\beta)$ are not used. [^1] If the search for pivots proceeded along columns rather than along rows, i.e. in the order last column top to bottom, then second last column top to bottom, etc., the additional restriction would be on the left rather than on the right.
@@
$T^{Right}\left(\beta\right)=\left\{(i,j):i\in Im(\beta),j<i,j\notin\left\{\beta(1),\dots,\beta\left(\beta^{-1}(i)-1\right)\right\}\right\}\subset Pairs\left(n\right)$ (56)
Note that these sets are transitive and behave naturally under extensions: for $\alpha^{\prime}$ an extension of $\alpha$, $T^{Left}\left(\alpha\right)\subset T^{Left}\left(\alpha^{\prime}\right)$, and for $\beta^{\prime}$ an extension of $\beta$, $T^{Right}\left(\beta\right)\subset T^{Right}\left(\beta^{\prime}\right)$.
@@
$(r,\alpha,\beta,L,R)$ can be computed from $A$ in time $O(mnr)$ by Gaussian elimination. [^2] Assuming $L,R$ are output as lists of off-diagonal non-zero positions; otherwise, outputting the full $L,R$ would take time $O(\max(m^{2},n^{2}))$.
Proof.
Encode the evolution of Gaussian elimination in a sequence: take $A_{0}=A$, and for $i\geq 1$ take
  1. 1.
    $(\alpha(i),\beta(i))$ to be the position of the $i$-th pivot in the left and down order.
  2. 2.
    $u_{i}=A_{i-1}e_{n,\beta(i)}-e_{m,\alpha(i)}$ to be the column of $A_{i-1}$ below the $i$-th pivot.
  3. 3.
    $v_{i}^{\top}=e_{m,\alpha(i)}^{\top}A_{i-1}-e_{n,\beta(i)}^{\top}$ to be the row of $A_{i-1}$ to the left of the $i$-th pivot.
  4. 4.
    $A_{i}=G_{m,u_{i},\alpha(i)}A_{i-1}G_{n,\beta(i),v_{i}^{\top}}$ to be the remaining matrix after row $\alpha(i)$ and column $\beta(i)$ have been cleared.
Induction shows that for each $i$, rows $1,\dots,\alpha(i)$ and columns $\beta(1),\dots,\beta(i)$ of $A_{i}$ match the corresponding rows and columns of $\sum_{k=1}^{i}e_{m,\alpha(k)}e_{n,\beta(k)}^{\top}$. Then, the sequence must terminate; let $A_{r}=\Pi(\alpha,\beta)$ be the last term. Then, $A=L\Pi(\alpha,\beta)R$ with
$L=G_{m,u_{1},\alpha(1)}G_{m,u_{2},\alpha(2)}\dots G_{m,u_{r},\alpha(r)}\in\mathcal{L}_{m}\left(T^{Left}\left(\alpha\right)\right)$ (58)
and
$R=G_{n,\beta{r},v_{r}^{\top}}G_{n,\beta(r-1),v_{r-1}^{\top}}\dots G_{n,\beta(1),v_{1}^{\top}}\in\mathcal{L}_{n}\left(T^{Right}\left(\beta\right)\right)$ (59)
Now, consider the complexity of the algorithm. For each $i$, $(\alpha(i),\beta(i),u_{i},v_{i},A_{i})$ can be computed from $(A_{i-1},\alpha(i-1))$ in time $O(mn)$. Then, the time to reach step $r$ is $O(mnr)$.
It remains to consider the time to compute $L$ and $R$. It is shown below that the non-zero off-diagonal positions can be computed in time $O(mr^{2})$ for $L$ and in time $O(nr^{2})$ for $R$.
Consider the sequence $L_{1}=G_{m,u_{1},\alpha(1)}$ and $L_{i}=L_{i-1}G_{m,u_{i},\alpha(i)}$ for $i=2,\dots,r$. Let $\tilde{L}_{i}$ be such that $L_{i}=I_{m}+\tilde{L}_{i}$. Then,
$\tilde{L}_{i}=\tilde{L}_{i-1}+u_{i}e_{m,\alpha(i)}^{\top}+\tilde{L}_{i-1}u_{i}e_{m,\alpha(i)}^{\top}$ (60)
Note that $\tilde{L}_{i-1},\tilde{L}_{i}$ are supported on some subsets of $T^{Left}\left(\alpha\right)$, so they have at most $r$ non-zero columns. Then, $\tilde{L}_{i}$ can be computed from $(\tilde{L}_{i-1},u_{i},\alpha(i))$ in time $O(mr)$, so $\tilde{L}=\tilde{L}_{r}$ can be computed from $(u_{1},\dots,u_{r},\alpha(1),\dots,\alpha(r))$ in time $O(mr^{2})$.
Similarly, consider the sequence $R_{1}=G_{n,\beta(1),v_{1}^{\top}}$ and $R_{i}=G_{n,\beta(i),v_{i}^{\top}}R_{i-1}$ for $i=2,\dots,r$. Let $\tilde{R}_{i}$ be such that $R_{i}=I_{n}+\tilde{R}_{i}$. Then,
$\tilde{R}_{i}=\tilde{R}_{i-1}+e_{n,\beta(i)}v_{i}^{\top}+e_{n,\beta(i)}v_{i}^{\top}\tilde{R}_{i-1}$ (61)
Note that $\tilde{R}_{i-1},\tilde{R}_{i}$ are supported on some subsets of $T^{Right}\left(\beta\right)$, so they have at most $r$ non-zero rows. Then, $\tilde{R}_{i}$ can be computed from $(\tilde{R}_{i-1},\beta(i),v_{i}^{\top})$ in time $O(nr)$, so $\tilde{R}=\tilde{R}_{r}$ can be computed from $(\beta(1),\dots,\beta(r),v_{1}^{\top},\dots,v_{r}^{\top})$ in time $O(nr^{2})$. ∎
@@
Take any $k\in[m],l\in[n]$, and divide the matrices into blocks with (k,mk)(k,m-k) and (l,nl)(l,n-l) rows/columns:
@@
Case 1: $i\leq j$. Then, $e_{n,i}^{\top}Re_{n,j}=e_{n,i}^{\top}e_{n,j}$ because $R$ is lower triangular.
Case 2: $i>j$, $i\notin Im(\beta)$. Then, $e_{n,i}^{\top}Re_{n,j}=0$ because $R\in\mathcal{L}_{n}\left(T^{Right}\left(\beta\right)\right)$.
Case 3: $i>j$, $\exists k:i=\beta(k)$, $j\notin Im(\beta)$. Then,
@@
Case 4: $i>j$, $\exists k,l:i=\beta(k),j=\beta(l),l<k$. Then, $e_{n,i}^{\top}Re_{n,j}=0$ because $R\in\mathcal{L}_{n}\left(T^{Right}\left(\beta\right)\right)$.
Case 5: $i>j$, $\exists k,l:i=\beta(k),j=\beta(l),k<l$. Then,
@@
Let $\alpha:[r]\rightarrow[m]$ be increasing, and let $L$ be in $\mathcal{L}_{m}\left(T^{Left}\left(\alpha\right)\right)$. If there exist $n$ and injective $\beta:[r]\rightarrow[n]$ such that $L\Pi(\alpha,\beta)=\Pi(\alpha,\beta)$, then $L=I$.
@@ 4.2 Canonical form for $m\times 2n$ stabilizer parity check matrices
$T^{moves}\left(\beta\right)=\left\{(i,j):i\in Im(\beta),j<i,q_{n}(j)\notin\left\{q_{n}(\beta(1)),\dots,q_{n}(\beta(\beta^{-1}(i)-1))\right\}\right\}$ (70)
@@
Case 1: $(i,j),(j,k)\in T^{moves}\left(\beta\right)$. Then, $i,j\in Im(\beta)$ with $\beta^{-1}(j)>\beta^{-1}(i)$. Moreover, $i>j>k$ and
@@
Case 2: $(i,j),(j,k)\in T^{rev}\left(\beta\right)$. Then, $(2n+1-k,2n+1-j),(2n+1-j,2n+1-i)\in T^{moves}\left(\beta\right)$, so $(2n+1-k,2n+1-i)\in T^{moves}\left(\beta\right)$, so $(i,k)\in T^{rev}\left(\beta\right)$.
Case 3: $(i,j)\in T^{moves}\left(\beta\right),(j,k)\in T^{rev}\left(\beta\right)$. Then $i>j>k$ and $i,2n+1-k\in Im(\beta)$. Consider subcases:
@@
Case 4: $(i,j)\in T^{rev}\left(\beta\right),(j,k)\in T^{moves}\left(\beta\right)$. Then, $j,2n+1-j\in Im(\beta)$. This is a contradiction, because $\beta$ is qubit injective. This case does not occur.
@@
Parts 2 and 3: Since $A$ is $(\alpha,\beta)$ reduced, and since $A\Omega_{2n}A^{\top}=0$, the following holds for the columns of $A$:
@@
Part 5: Computation of $(\alpha^{\prime},\beta^{\prime},u,v)$ can clearly be done in time $O(mn)$. For the computation of $A^{\prime}$, use the sparsity of $G_{m,u,\alpha^{\prime}(r+1)},S_{2n,\beta^{\prime}(r+1),v^{\top}}$. For any $x\in\mathbb{F}_{2}^{m}$, $G_{m,u,\alpha^{\prime}(r+1)}x$ can be computed in time $O(m)$; since $A$ has $2n$ columns, this gives time $O(mn)$ to compute $G_{m,u,\alpha^{\prime}(r+1)}A$. Similarly, for any $y\in\mathbb{F}_{2}^{2n}$, $y^{\top}S_{2n,\beta^{\prime}(r+1),v^{\top}}$ can be computed in time $O(n)$; for $m$ rows, this gives time $O(mn)$ to compute $G_{m,u,\alpha^{\prime}(r+1)}AS_{2n,\beta^{\prime}(r+1),v^{\top}}$. ∎
@@
Case 1: $(i,j)\in T^{moves}\left(\beta\right)$. Then, $i=\beta(k)$ for some $k\in[r]$ and
@@
Case 2: $(i,j)\in T^{rev}\left(\beta\right)$. Then, $(2n+1-j,2n+1-i)\in T^{moves}\left(\beta\right)$ and $2n+1-j=\beta(k)$ for some $k\in[r]$. Apply the identity $R^{\top}=\Omega_{2n}R^{-1}\Omega_{2n}$, which holds for any element of $\mathcal{SP}_{2n}$:
@@
    Moreover, the algorithm runs in time $O(mnr)$ (assuming $L,R$ are output as lists of off-diagonal non-zero positions; otherwise, outputting the full $L,R$ would take time $O(\max(m^{2},n^{2}))$).
@@
Part 1 follows by repeated application of Lemma 8. The elimination step in Lemma 8 needs to be performed $r$ times, so the total time is $O(mnr)$. It remains to consider the time to compute $L$ and $R$. It was shown in the proof of theorem 3 that the non-zero off-diagonal entries of $L\in\mathcal{L}_{n}\left(T^{Left}\left(\alpha\right)\right)$ can be computed in time $O(mr^{2})$. Next, consider the computation of
$R=S_{2n,\beta(r),v_{r}^{\top}}S_{2n,\beta(r-1),v_{r-1}^{\top}}\dots S_{2n,\beta(1),v_{1}^{\top}}\in\mathcal{B}_{2n}\left(T^{tcr}\left(\beta\right)\right)$ (86)
where $v_{1}^{\top},\dots,v_{r}^{\top}$ is the sequence of vectors that arises from the $r$ applications of Lemma 8. Consider the sequence $R_{1}=S_{2n,\beta(1),v_{1}^{\top}}$ and $R_{i}=S_{2n,\beta(i),v_{i}^{\top}}R_{i-1}$ for $i=2,\dots,r$. Let $\tilde{R}_{i}$ be such that $R_{i}=I_{2n}+\tilde{R}_{i}$ and let $\tilde{S}_{i}$ be such that $S_{2n,\beta(i),v_{i}^{\top}}=I_{2n}+\tilde{S}_{i}$. Then,
$\tilde{R}_{i}=\tilde{R}_{i-1}+\tilde{S}_{i}+\tilde{S}_{i}\tilde{R}_{i-1}$ (87)
Note that $\tilde{R}_{i-1},\tilde{R}_{i}$ are supported on some subsets of $T^{tcr}\left(\beta\right)$, which in turn is a subset of $Pairs\left(2n\right)\cap((Im(\beta)\times[2n])\cup([2n]\times(2n+1-Im(\beta))))$. Note also that $\tilde{S}_{i}$ has non-zero entries only in row $\beta(i)$ and column $2n+1-\beta(i)$. Therefore,
Claim: $\tilde{R}_{i}$ can be computed from $(\tilde{R}_{i-1},\tilde{S}_{i})$ in time $O(nr)$.
Proof of claim: Partition $T^{tcr}\left(\beta\right)$ into three sets:
$T_{1}$ $=\left\{(k,l)\in T^{tcr}\left(\beta\right):k\neq\beta(i)\right\}$ (88)
$T_{2}$ $=\left\{(k,l)\in T^{tcr}\left(\beta\right):k=\beta(i),l\notin(2n+1-Im(\beta))\right\}$ (89)
$T_{3}$ $=\left\{(k,l)\in T^{tcr}\left(\beta\right):k=\beta(i),l\in(2n+1-Im(\beta)\right\}$ (90)
and note that $|T_{1}|=O(nr)$, $|T_{2}|=O(n)$, $|T_{3}|=O(r)$. For each $(k,l)\in T_{1}$,
$e_{2n,k}^{\top}\tilde{R}_{i}e_{2n,l}\\ =e_{2n,k}^{\top}\tilde{R}_{i-1}e_{2n,l}+e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,l}+e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,2n+1-\beta(i)}e_{2n,2n+1-\beta(i)}^{\top}\tilde{R}_{i-1}e_{2n,l}$ (91)
can be computed in time $O(1)$. For each $(k,l)\in T_{2}$,
$e_{2n,k}^{\top}\tilde{R}_{i}e_{2n,l}=e_{2n,k}^{\top}\tilde{R}_{i-1}e_{2n,l}+e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,l}+\sum_{j\in Im(\beta)}e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,j}e_{2n,j}^{\top}\tilde{R}_{i-1}e_{2n,l}$ (92)
can be computed in time $O(r)$. For each $(k,l)\in T_{3}$,
$e_{2n,k}^{\top}\tilde{R}_{i}e_{2n,l}=e_{2n,k}^{\top}\tilde{R}_{i-1}e_{2n,l}+e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,l}+\sum_{j\in[2n]}e_{2n,k}^{\top}\tilde{S}_{i}e_{2n,j}e_{2n,j}^{\top}\tilde{R}_{i-1}e_{2n,l}$ (93)
can be computed in time $O(n)$. This completes the proof of the claim.
Using the claim, deduce that the non-zero off-diagonal positions of $R$ can be computed from $(\beta(1),\dots,\beta(r),v_{1}^{\top},\dots,v_{r}^{\top})$ in time $O(nr^{2})$. This completes the proof of Part 1.